OpenStudy (anonymous):

(3^8)/ (3^3) I know it is 243 but how do you get that?

OpenStudy (anonymous):

I was told that when you are dividing like bases with exponents you just simplify the base as normal and subtract the exponents. That is no t a choice however

OpenStudy (anonymous):

3^8/3^3=3^5=3*3*3*3*3=243

OpenStudy (anonymous):

Oh! Thank you!